OpenStudy (anonymous):

A stretch of river has straight parallel banks 120m apart. A man rowing at 5m/s relative to the water and heading directly across the river reaches the other bank 54m downstream. Find the speed of the current and the direction in which the man would have to row in order to cross the river at right angles to the banks. .

OpenStudy (anonymous):

 now first we have to find the speed of river so we will find angle first costheta=120/5 =24degree

OpenStudy (anonymous):

now for finding river speed sintheta=Mr/5 5x5/120=Mr 0.11m/sec

OpenStudy (anonymous):

go on...

OpenStudy (anonymous):

now other part